Following question:
-Calulated molarity of diluted vinegar solution
-Molarity of Commercial Vinegar (multiply by the dilution factor
-Percent (mass/volume) of acetic acid in commefcial vinegar
Data and Calculations Molarity of standardized sodium hydroxide solution Dilution factor for vinegar solution prepared in step 4 i 20 Volume of diluted vinegar titrated in each run Titrations: Run 3 Run 2 Run 1 G (o .00 ml Final Buret 22.20 Volume (mL) Initial Buret Volume (mL) 22.20 WL Volume NaOH 22.3 Delivered (mL) Average Volume of NaOH (mL) Calculated molarity of diluted vinegar solution:
media%2F93c%2F93ce06f8-a236-4345-883e-1f
Answer #1

Given

Volume of acetic acid V1 = 25 ml

Volume of NaOH V2 = 22.27 ml

Molarity of NaOH M2 = 0.09766 mol/L

V1 * M1 = V2 * M2

25 ml * M1 = 22.27 ml * 0.09766 mol/L

M1 = 22.27 * 0.09766 / 25 = 0.087 mol/L

Molarity of diluted vinegar solution = 0.087 mol/L

Molarity of commercial vinegar = 0.087 mol/L * dilution factor = 0.087 * 10 = 0.87 mol/L

Mol. wt of Vinegar = 60 g/mol

percent (mass/volume) = molarity * mol. wt * 100 % = 0.87 mol/L * 60 g/mol * 100 %

= (0.87 mol/ 1000 ml) * 60 g/mol * 100 % = 5.22 % answer